If the MC = q, which is the marginal abatement costs, and MB= 10-q, which is the marginal benefits of abatement, what are the welfare gains from socially optimal regulations?
Answer #1

In social optimal, MB = MC.

10 - q = q

2q = 10

q = 5

P = 5

When q = 0, MB = 10

Consumer surplus (CS) = (1/2) x (10 - 5) x 5 = 2.5 x 5 = 12.5

When q = 0, MC = 0

Producer surplus (PS) = (1/2) x (5 - 0) x 5 = 2.5 x 5 = 12.5

Total welfare gain = CS + PS = 12.5 + 12.5 = 25